Grade 12Physical Chemistry

A certain particle carries 2.5 × 10–16C of static electric charge. Calculate the number of electrons present in it

sudhanshu
12 Years agoGrade 12

1 Answer

Gaurav
11 Years ago
Hello Student
Let the number of electrons present in ion A3+ be x



∴ Number of neutrons in it =x+ 30.4% ofx= 1.304x

Since the ion is tripositive,

⇒ Number of electrons in neutral atom =x+ 3

∴ Number of protons in neutral atom =x+ 3

Given,

Mass number of the ion = 56

(x+3)+(1.304x) = 56

2.304x = 53

x = 23

∴ Number of protons =x+ 3 = 23 + 3 = 26

∴ The symbol of the ion Fe3+